OpenStudy (anonymous):

CALC (Will medal) Finding the derivative of the function below... I know I can use second fundamental theorem but how?

OpenStudy (anonymous):

\[-x \int\limits_{2}^{x}f(t)dt\]

OpenStudy (anonymous):

Would it just be \[-\int\limits_{2}^{x}f(x)\]

OpenStudy (anonymous):

Making sure, so this is your problem correct? \[-x \int\limits_{2}^{x} f(t) dt\]

OpenStudy (anonymous):

Yes

OpenStudy (zale101):

What is the f(t) equation? Is it given?

OpenStudy (anonymous):

Not really, the only equation given is y = f(x)

OpenStudy (anonymous):

*only other equation

OpenStudy (anonymous):

You have to use fundamental theorem of calc 1, I'm trying to remember so \[g(x) = -x \int\limits\limits_{2}^{x} f(t) dt \] then \[g'(x) = f(x)\]

OpenStudy (anonymous):

Oh, that's right! Completely forgot to remove integral. Thank you for your help!

OpenStudy (anonymous):

Thinking further... wouldn't it be -1 + f(x) due to the derivative of -x being -1?